a. Draw a consumer budget constraint and indifference curves for Pepsi and Pizza. Find out the optimal consumption choice when consumer has income of $2000 and the price of Pepsi is $5 per bottle and the price of per pizza is $10.What is the marginal rate of substitution at this optimum? (6) b. Piyush spends his entire income between food grains and books (both of which are normal goods) Pandemic causes a large increase in the prices of food grains in India.


i. Show the effect of Pandemic on Piyush’s budget constraint.


ii. Show the effect of Pandemic on Piyush’s optimal consumption bundle assuming that substitution effect outweighs the income effect for books.


iii. Show the effect of Pandemic on Piyush’s optimal consumption bundle assuming that income effect outweighs the substitution effect for books.


1
Expert's answer
2021-03-29T12:07:05-0400


(i) Books and grains are normal goods meaning an increase in price of grains will lead to increased quantity demanded for books than grains.

Piyush's budget constraint before pandemic was originally

"(PF\times QF)+(PB\times QB)=Y(PF\times QF)+(PB\times QB)=Y"

PF=price of grains

QB=quantity of books consumed.

QF=quantity of food grains consumed

PB=price of books

After increase in price of food grains, the price of food grains increases from PF to PF1 and the quantity decreases from QF to QF1. The new budget constraint will be

"(PF1\times QF1)+(PB\times QB)=Y(PF1\times QF1)+(PB\times QB)=Y"

(ii) substitution effect will increase the quantity demanded for books from QB to QB1)

"(PF1\times QF1)+(PB\times QB1)=Y(PF1\times QF1)+(PB\times QB1)=Y"

(iii) Income effect increases the income from Y to Y1

"(PF1\times QF1)+(PB\times QB1)=Y1(PF1\times QF1)+(PB\times QB1)=Y1"